# Constacyclic codes with best-known parameters

### v1

@@ Abstract
In this paper, we construct several infinite families of $q$-ary constacyclic codes over a finite field $\mathbb{F}_{q}$ with length $n$, dimension around $n/2$, and minimum distance at least $cn/\log_{q}n$ for some positive constant $c$. They contain many constacyclic codes with optimal, or almost-optimal, or best-known parameters. We also consider various forms of the length $n$.
@@ 1.1 Background
Let $\mathbb{F}_{q}$ be the finite field with $q$ elements and $\mathbb{F}_{q}^{*}=\mathbb{F}_{q}\setminus\{0\}$. A $q$-ary $[n,k,d]$ linear code $\mathcal{C}$ is a $k$-dimensional linear subspace of $\mathbb{F}_{q}^{n}$ with minimum distance $d$. Here, $n$ is called the length of $\mathcal{C}$. The linear code $\mathcal{C}$ is said to be distance optimal if there is no $q$-ary $n,k,\geq d+1]$ linear code. In addition, $\mathcal{C}$ is said to be distance almost-optimal if there is a $q$-ary $[n,k,d+1]$ linear code but there is no $q$-ary $n,k,\geq d+2]$ linear code. Moreover, $\mathcal{C}$ is said to be distance best-known if $d$ is the maximal minimum distance of all known $q$-ary linear codes with length $n$ and dimension $k$.
Let $\lambda\in\mathbb{F}_{q}^{*}$. The linear code $\mathcal{C}$ is called a $\lambda$-constacyclic code if $(\lambda c_{n-1},c_{0},\ldots,c_{n-2})\in\mathcal{C}$ for any $(c_{0},c_{1},\ldots,c_{n-1})\in\mathcal{C}$. Define the map
@@
It is well-known that every ideal of the residue class ring $\mathbb{F}_{q}[x]/\langle x^{n}-\lambda\rangle$ is principal. Then, the linear code $\mathcal{C}$ is called a $\lambda$-constacyclic code if and only if $\phi(\mathcal{C})$ is an ideal of the ring $\mathbb{F}_{q}[x]/\langle x^{n}-\lambda\rangle$. So, we identify $\mathcal{C}$ with $\phi(\mathcal{C})$ for any $\lambda$-constacyclic code $\mathcal{C}$. Let $\mathcal{C}=\langle g(x)\rangle$ be a $q$-ary $\lambda$-constacyclic code of length $n$, where $g(x)$ is monic and has the smallest degree among all the non-zero codewords of $\mathcal{C}$. Then, $g(x)$ is a factor of $x^{n}-\lambda$ and is called the generator polynomial of $\mathcal{C}$.
@@
Constacyclic codes have attracted extensive attentions due to their performance and applications. For example, several new constacyclic codes that improve the minimum distance of currently best known linear codes were found in [1, 2], several infinite classes of MDS constacyclic codes were constructed in [14, 18, 36], and several infinite families of distance-optimal constacyclic codes were constructed in [12, 42]. In addition, constacyclic codes have important applications in constructing other kinds of codes, such as locally repairable codes [3], symbol-pair codes [4], and quantum codes [5, 17]. We refer to [30] for more details and references about constacyclic codes.
@@ 1.2 Related works
It is a long-standing open problem in coding theory that whether there exists an asymptotically good sequence of cyclic codes (see [25]). Therefore, it is interesting to construct infinite families of cyclic, negacyclic or constacyclic codes whose dimensions and minimum distances are both large.
In recent years, researchers pay much attention in constructing infinite families of such codes with parameters $[n,k,d]$ such that $k$ is around $n/2$ and $d$ is at least $c\sqrt{n}$ for some positive constant $c$ depending only on $q$ (this lower bound is called the square-root-like lower bound).
Some binary quadratic residue codes with parameters $[n,(n+1)/2,d]$ and their even-weight subcodes with parameters [n,(n1)/2,d+1][n,(n-1)/2,d+1], where $d\geq\sqrt{n}$ and $n$ is a prime with $n\equiv\pm 1\pmod{8}$, were constructed in [16, Section 6.6]. Two infinite families of binary cyclic codes with length $2^{m}-1$ and dimension near $2^{m-1}$ and with a square-root-like lower bound on minimum distance were constructed in [34]. Recently, the work in [34] was extended to the 4-ary case in [28] and the $2^{s}$-ary case in [19]. We refer to [20, 21, 22, 29, 31] for more infinite families of binary cyclic codes with a square-root-like lower bound, [10, 22] for infinite families of ternary cyclic codes with a square-root-like lower bound, and [6, 7, 33, 40, 41] for infinite families of $q$-ary cyclic codes with a square-root-like lower bound. In addition, several infinite families of $q$-ary negacyclic codes and constacyclic codes were constructed in [11, 30, 38] with a square-root-like lower bound.
@@
Motivated by [32], several works have been done recently for constructing more $q$-ary linear codes with a lower bound like $cn/\log_{q}n$ on their minimum distances. We summarize them as follows:
  • In [8], one infinite family of $q$-ary negacyclic $n,(n+1)/2,\geq cn/\log_{q}n]$ codes with $q\equiv 3\pmod{4}$ and $n=\frac{q^{p}-1}{q-1}$ was constructed, and one infinite family of $q$-ary negacyclic $n,n/2,\geq cn/\log_{q}n]$ codes with $q\equiv 1\pmod{4}$ and $n=\frac{q^{p}-1}{2}$ was also constructed.
@@
  • In [20], two infinite families of binary $n,(n+1)/2,\geq cn/\log_{2}n]$ cyclic codes with $n=2^{p^{2}}-1$ or $2^{p_{1}p_{2}}-1$ were construced in [20], where $p_{1}$ and $p_{2}$ are two distinct primes.
  • In [23], an improvement of the lower bound on minimum distance in [32] was given, and an infinite family of ternary $n,(n+2)/2,\geq cn/\log_{3}n]$ cyclic codes was constructed, where $n=3^{p}-1$.
@@
Most recently, an infinite family of binary BCH codes was constructed in [39] with length $n=(2^{p_{1}}-1)\cdots(2^{p_{s}}-1)$, dimension at least $(n+1)/2$, and minimum distance at least $\lceil(n-1)/(p_{1}\cdots p_{s})\rceil$, where $p_{1},\ldots,p_{s}$ are distinct primes. We also refer to [26] for BCH codes with minimum distances propotional to code lengths.
@@ 1.3 Our contributions
In this paper, motivated by [32] and generalizing the constructions in [8, 9, 20, 23, 31, 32], we propose the following general construction for $q$-ary constacyclic codes: first we classify the cyclotomic cosets according to their sizes (that is, two cosets are in the same class if and only if they have the same size), and then we take the defining set as the union of the first half of cosets in each class (we will use the ceiling function and the floor function respectively to achieve “the first half”).
In order to obtain good lower bounds for minimum distance, in this paper we focus on the case when there are exactly two kinds of cyclotomic cosets. For the minimum distance, if $N$ is the number of cyclotomic cosets with maximum size, by construction one may guess that the first $\lfloor N/2\rfloor$ consecutive elements are in the defining set. However, this does not always hold.
@@
  • We establish the lower bound $\lfloor\frac{qN}{2(q-1)}\rfloor$ on minimum distances for $q$-ary constacyclic codes constructed by using the ceiling function, and the same lower bound in many situations for those codes constructed by using the floor function (see Theorem 3.1). This improves the lower bounds on minimum distance in [8, Theorems 10, 15 and 16] for some infinite families of $q$-ary negacyclic codes mentioned above. Moreover, for some special cases, we obtain better lower bounds on minimum distances (see Theorem 3.3).
  • We consider various forms of the length $n$: $n=p$, or $\frac{q^{p}-1}{rs}$, or $\Phi_{p^{b}}(q)$, or $\Phi_{p_{1}p_{2}}(q)$, where $p$ is a prime, $r$ is the multiplicative order of $\lambda$, $s$ and $b$ are two positive integers, $p_{1}$ and $p_{2}$ are two distinct primes, and $\Phi_{m}(x)$ is the $m$-th cyclotomic polynomial for any integer $m\geq 1$. We obtain many constacyclic codes with optimal, or almost-optimal, or best-known parameters (according to [15]).
The rest of this paper is organized as follows. In Section 2, we present some preliminary results. In Section 3, we present some general constructions for infinite families of constacyclic codes. In Section 4, we construct two infinite families of constacyclic codes with prime length. In Section 5, we construct two infinite families of constacyclic codes with length $n=\frac{q^{p}-1}{rs}$. In Section 6, we construct several infinite families of constacyclic codes with length $n=\Phi_{p^{b}}(q)$. In Section 7, we construct two infinite families of constacyclic codes with length $n=\Phi_{p_{1}p_{2}}(q)$.
@@ 2.1 Cyclotomic cosets
where $l_{i}$ is the least positive integer such that $iq^{l_{i}}\equiv i\mod nr$, that is,
@@
The smallest integer in $C_{i}^{(q,nr)}$ is called the coset leader of $C_{i}^{(q,nr)}$. Let $\Gamma_{(q,n,r)}$ be the set of all the coset leaders,
@@
$M_{\beta^{i}}(x)=\prod_{j\in C_{i}^{(q,nr)}}(x-\beta^{j})\in\mathbb{F}_{q}[x].$
@@
is called the defining set of $\mathcal{C}$ with respect to $\beta$. In addition, the Bose distance of $\mathcal{C}$ is defined to be the largest number of consecutive elements from $Z_{n,r}$ and contained in $T(\mathcal{C})$ (see [13] for some work about Bose distance).
For any positive integer $m$ with $\gcd(q,m)=1$, let $\mathrm{ord}_{m}(q)$ be the the multiplicative order of $q$ modulo $m$. The result in the following lemma is somehow well-known (for instance, see [35, page 3]).
@@
For any positive integer $l$, let $N_{l}^{(q,nr)}$ be the number of $q$-cyclotomic cosets modulo $nr$ whose coset leaders are in $\Gamma_{q,n,r}^{(1)}$ and sizes are equal to $l$, that is,
$N_{l}^{(q,nr)}=\Big|\big\{C_{i}^{(q,nr)}:\,|C_{i}^{(q,nr)}|=l,i\in Z_{n,r}\big\}\Big|.$
By Lemma 2.1, if $l\nmid\mathrm{ord}_{nr}(q)$, then $N_{l}^{(q,nr)}=0$. In addition, $N_{l}^{(q,nr)}=0$ if $l\neq l_{i}$ for any $i\in Z_{n,r}$.
@@
$N_{l}=N_{l}^{(q,nr)},\quad l\geq 1.$
We recall two lemmas in [35, Lemmas 4 and 5], which are useful for determining $N_{l}$.
@@
From Lemma 2.3, we can get a formula about the number of elements $i\in Z_{n,r}$ with $l_{i}$ dividing a given integer.
@@
Proof.
By Lemma 2.3, we have
$lN_{l}=\sum_{j\mid l}\mu(l/j)\sigma(j)\gcd(n,(q^{j}-1)/r).$
Using the Möbius inversion formula, we directly have
$\sum_{j\mid l}jN_{j}=\sigma(l)\gcd(n,(q^{l}-1)/r),$
which gives the desired result. ∎
@@ 2.2 Coset leaders in $Z_{n,r}$
Now, for any positive integer $l$, if $N_{l}^{(q,nr)}>0$, then in $Z_{n,r}$, we list all the coset leaders (in ascending order) corresponding to those $q$-cyclotomic cosets with $l$ elements modulo $nr$, and we denote the $i$-th one by $\delta_{i}^{(q,nr,l)}$ for $1\leq i\leq N_{l}^{(q,nr)}$.
@@
$N_{l}=N_{l}^{(q,nr)},\quad l\geq 1,$
$\delta_{i}^{(l)}=\delta_{i}^{(q,nr,l)},\quad 1\leq i\leq N_{l}.$
We want to get some estimates about the coset leaders $\delta_{i}^{(l)}$ for $1\leq i\leq N_{l}$. For this, we need a preparation.
@@
Assume that the $i$-th element not disivible by $q$ in the set $S_{a,r}$ is a+(j1)ra+(j-1)r for some integer $j$. Then, we have either $i=j-\lfloor j/q\rfloor-1$, or $i=j-\lfloor j/q\rfloor$. If $i=j-\lfloor j/q\rfloor-1$, then we have
$i\leq j-(j/q-(q-1)/q)-1,$
which gives
$j\geq\frac{qi}{q-1}+\frac{1}{q-1}.$
If $i=j-\lfloor j/q\rfloor$, then we have
$i\leq j-(j/q-(q-1)/q),$
which implies
$j\geq\frac{qi}{q-1}-1.$
Hence, noticing that $j$ is an integer, we always have
@@
First,by Lemma 2.2 we have
$\frac{nr}{\gcd(nr,q^{l}-1)}\,\Big|\,\delta_{i}^{(l)}.$
For simplicity, we denote $t=nr/\gcd(nr,q^{l}-1)$, then $\delta_{i}^{(l)}$ is a multiple of $t$. Since $r\mid q-1$, we have $t\mid n$.
@@
$S(t)=\{at,at+rt,at+2rt,\ldots,at+jrt\},$
where $j$ is the largest integer such that $at+jrt\leq 1+(n-1)r$. Clearly, we have
$S(t)=\{at,(a+r)t,(a+2r)t,\ldots,(a+jr)t\}.$
Hence, for our purpose it suffices to estimate the $i$-th element in $S(t)$ (listed in ascending order) which is not disivible by $q$.
@@
$\begin{split}\delta_{i}^{(l)}&\geq\left(a+(\lceil qi/(q-1)\rceil-2)r\right)t\\ &\geq\frac{nr(1+(\lceil qi/(q-1)\rceil-2)r)}{\gcd(nr,q^{l}-1)}.\end{split}$
@@
The following lemma is a generalization of [32, Lemma 9] for determining some elements which are not coset leaders.
@@
Moreover, all these $q^{i}+qt+j$ are pairwise distinct.
@@
$\begin{split}q^{i}+qt+j&<q^{i}+q\cdot\frac{q^{i}-(q^{2}-1)-1}{q^{3}-q}+q-1\\ &<2q^{i}\leq q^{i+1}.\end{split}$ (2.1)

### v2

@@ Abstract
In this paper, we construct several infinite families of $q$-ary constacyclic codes over a finite field $\mathbb{F}_{q}$ with length $n$, dimension around $n/2$, and minimum distance at least $cn/\log_{q}n$ for some positive constant $c$. They contain many constacyclic codes with optimal, or almost-optimal, or best-known parameters. We also consider constacyclic codes of various lengths.
@@ 1.1 Background
Let $\mathbb{F}_{q}$ be the finite field with $q$ elements and $\mathbb{F}_{q}^{*}=\mathbb{F}_{q}\setminus\{0\}$. A $q$-ary $[n,k,d]$ linear code $\mathcal{C}$ is a $k$-dimensional linear subspace of $\mathbb{F}_{q}^{n}$ with minimum distance $d$. Here, $n$ is called the length of $\mathcal{C}$. The code $\mathcal{C}$ is said to be distance-optimal (or, optimal, for short) if there is no $q$-ary $n,k,\geq d+1]$ linear code. In addition, $\mathcal{C}$ is said to be distance almost-optimal (or, almost-optimal, for short) if there is a distance-optimal $q$-ary $[n,k,d+1]$ linear code. Moreover, $\mathcal{C}$ is said to be best-known if $d$ is the maximal minimum distance of all known $q$-ary linear codes with length $n$ and dimension $k$.
Let $\lambda\in\mathbb{F}_{q}^{*}$. The code $\mathcal{C}$ is called a $\lambda$-constacyclic code if $(\lambda c_{n-1},c_{0},\ldots,c_{n-2})\in\mathcal{C}$ for any $(c_{0},c_{1},\ldots,c_{n-1})\in\mathcal{C}$. Define the map
@@
It is well-known that every ideal of the residue class ring $\mathbb{F}_{q}[x]/\langle x^{n}-\lambda\rangle$ is principal. Then, the code $\mathcal{C}$ is called a $\lambda$-constacyclic code if and only if $\phi(\mathcal{C})$ is an ideal of the ring $\mathbb{F}_{q}[x]/\langle x^{n}-\lambda\rangle$. Moreover, we identify $\mathcal{C}$ with $\phi(\mathcal{C})$, and then we can write $\mathcal{C}=\langle g(x)\rangle$, where $g(x)$ is monic and has the smallest degree among all the non-zero codewords of $\mathcal{C}$. Here, $g(x)$ is in fact a factor of $x^{n}-\lambda$ and is called the generator polynomial of $\mathcal{C}$.
@@
Constacyclic codes have attracted extensive attentions due to their performance and applications. For example, several new constacyclic codes that improve the minimum distance of currently best known linear codes were found in [1, 2], and several infinite classes of MDS constacyclic codes were constructed in [13, 22, 43]. In addition, constacyclic codes have important applications in constructing other kinds of codes, such as locally repairable codes [4], symbol-pair codes [5], and quantum codes [6, 21]. We refer to [37] for more details and references about constacyclic codes.
@@ 1.2 Related works
Constructing cyclic codes or more general constacyclic codes with optimal or best-known parameters is interesting and important. Many optimal or best-known codes over small fields with short lengths are cyclic codes; see [18]. Some infinite families of optimal ternary cyclic codes with length $3^{m}-1$ and minimum distance four were constructed in [16]. Several other infinite families of optimal ternary cyclic codes with length $3^{m}-1$ and minimum distance four or five were constructed in [24]. Some infinite families of optimal ternary negacyclic codes with minimum distance four or five were constructed in [36]. Infinitely many families of optimal binary cyclic codes with minimum distance six and an infinite family of quaternary cyclic codes with minimum distance four were constructed in [9]. The first infinite family of binary optimal BCH codes with minimum distance 8 was constructed in [47]. Three infinite families of optimal repeated-root cyclic codes with minimum distance three or four were constructed in [10]. We refer to [9, 10, 14, 27, 36, 45, 47, 49, 51] and the references therein for more optimal or best-known codes constructed as cyclic, negacyclic, or quasi-cyclic codes.
On the other hand, it is a long-standing open problem in coding theory that whether there exists an asymptotically good sequence of cyclic codes (see [31]). So, it is interesting to construct infinite families of cyclic, negacyclic or constacyclic codes whose dimensions and minimum distances are both large.
In recent years, researchers have paid much attention in constructing infinite families of such codes with parameters $[n,k,d]$ such that $k$ is around $n/2$ and $d$ is at least $c\sqrt{n}$ for some positive constant $c$ depending only on $q$ (this lower bound is called the square-root-like lower bound).
Some binary quadratic residue codes with parameters $[n,(n+1)/2,d]$, where $d\geq\sqrt{n}$ and $n$ is a prime with $n\equiv\pm 1\pmod{8}$, were constructed in [20, Section 6.6]. Two infinite families of binary cyclic codes with length $2^{m}-1$ and dimension near $2^{m-1}$ and with a square-root-like lower bound on the minimum distance were constructed in [41]. Recently, the work in [41] was extended to the 4-ary case in [34] and the $2^{s}$-ary case in [23]. We refer to [25, 26, 28, 35, 38] for more infinite families of binary cyclic codes with a square-root-like lower bound, [11, 28] for infinite families of ternary cyclic codes with a square-root-like lower bound, and [8, 40, 48, 50] for infinite families of $q$-ary cyclic codes with a square-root-like lower bound. In addition, several infinite families of $q$-ary negacyclic codes and constacyclic codes were constructed in [12, 37, 46] with a square-root-like lower bound.
@@
Motivated by [39], several works have been done recently for constructing more $q$-ary cyclic or negacyclic codes with a lower bound like $cn/\log_{q}n$ on their minimum distances. We summarize them as follows:
  • In [9], one infinite family of $q$-ary $n,(n+1)/2,\geq cn/\log_{q}n]$ negacyclic codes with $q\equiv 3\pmod{4}$ and $n=\frac{q^{p}-1}{q-1}$ was constructed, and one infinite family of $q$-ary negacyclic $n,n/2,\geq cn/\log_{q}n]$ codes with $q\equiv 1\pmod{4}$ and $n=\frac{q^{p}-1}{2}$ was also constructed.
@@
  • In [25], two infinite families of binary $n,(n+1)/2,\geq cn/\log_{2}n]$ cyclic codes with $n=2^{p^{2}}-1$ or $2^{p_{1}p_{2}}-1$ were constructed, where $p_{1}$ and $p_{2}$ are two distinct odd primes.
  • In [29], an improvement of the lower bound on the minimum distance in [39] was given, and an infinite family of ternary $n,(n+2)/2,\geq cn/\log_{3}n]$ cyclic codes was constructed, where $n=3^{p}-1$.
@@
Most recently, an infinite family of binary BCH codes was constructed in [47] with length $n=(2^{p_{1}}-1)\cdots(2^{p_{s}}-1)$, dimension at least $(n+1)/2$, and minimum distance at least $\lceil(n-1)/(p_{1}\cdots p_{s})\rceil$, where $p_{1},\ldots,p_{s}$ are distinct primes. We also refer to [32] for BCH codes with minimum distances propotional to code lengths and refer to [17] for a recent survey on BCH codes.
@@ 1.3 Our contributions
In this paper, motivated by [39] and generalizing the constructions in [9, 10, 25, 29, 38, 39], we propose the following general construction for $q$-ary constacyclic codes: first we classify the cyclotomic cosets according to their sizes (that is, two cosets are in the same class if and only if they have the same size), and then we take the defining set as the union of the first half of cosets in each class (we sort the cosets in ascending order by their coset leaders and use the ceiling function or the floor function to achieve “the first half”).
In order to obtain good lower bounds for minimum distance, in this paper we focus on the case when there are exactly two kinds of cyclotomic cosets. Even in this case, the codes we construct are not always BCH constacyclic codes; see Example 3.3.
Let $N$ be the number of cyclotomic cosets with maximum size. In the above construction, one may guess that the first $\lfloor N/2\rfloor$ consecutive elements are in the defining set. However, this does not always hold; see Example 2.14.
@@
  • We establish the lower bound $\lfloor\frac{qN}{2(q-1)}\rfloor$ on minimum distances for $q$-ary constacyclic codes constructed by using the ceiling function, and the same lower bound in many situations for those codes constructed by using the floor function (see Theorem 3.1). This improves the lower bounds on minimum distances in [9, Theorems 10, 15 and 16] for some infinite families of $q$-ary negacyclic codes mentioned above. Moreover, for some special cases, we obtain better lower bounds on minimum distances (see Theorem 3.4).
  • We consider $q$-ary constacyclic codes of various lengths $n$: $\frac{q^{p}-1}{rs}$, or $\Phi_{p^{b}}(q)$, or $\Phi_{p_{1}p_{2}}(q)$, where $p$ is a prime, $r$ is the multiplicative order of $\lambda$, $s$ and $b$ are two positive integers, $p_{1}$ and $p_{2}$ are two distinct primes, and $\Phi_{m}(x)$ is the $m$-th cyclotomic polynomial for any integer $m\geq 1$. We obtain many constacyclic codes with optimal, or almost-optimal, or best-known parameters (according to [18]). We also present several codes which are better than BCH constacyclic codes.
The rest of this paper is organized as follows. In Section 2, we present some preliminary results. In Section 3, we present some general constructions for infinite families of constacyclic codes. In Section 4, we construct two infinite families of constacyclic codes with length $n=\frac{q^{p}-1}{rs}$ for any positive integers $r,s$ such that $r\mid q-1$ and $s\mid\frac{q^{p}-1}{r}$. In Section 5, we construct several infinite families of constacyclic codes with length $n=\Phi_{p^{b}}(q)$. In Section 6, we construct two infinite families of constacyclic codes with length $n=\Phi_{p_{1}p_{2}}(q)$.
@@ 2.1 Cyclotomic cosets
where $l_{i}$ is the least positive integer such that $iq^{l_{i}}\equiv i\pmod{nr}$, that is,
@@
For simplicity, in the sequel we denote $C_{i}=C_{i}^{(q,nr)}$.
The smallest integer in $C_{i}$ is called the coset leader of $C_{i}$. Let $\Gamma_{(q,n,r)}$ be the set of all the coset leaders,
@@
$M_{\beta^{i}}(x)=\prod_{j\in C_{i}}(x-\beta^{j})\in\mathbb{F}_{q}[x].$
@@
is called the defining set of $\mathcal{C}$ with respect to $\beta$. In addition, the Bose distance of $\mathcal{C}$ is defined to be one plus the largest number of consecutive elements from $Z_{n,r}$ and contained in $T(\mathcal{C})$ (see [15] for some work about the Bose distance of BCH codes).
For any positive integer $m$ with $\gcd(q,m)=1$, let $\mathrm{ord}_{m}(q)$ be the the multiplicative order of $q$ modulo $m$. The result in the following lemma is somehow well-known (for instance, see [42, page 3]).
@@
For any positive integer $l$, let $N_{l}^{(q,n,r)}$ be the number of $q$-cyclotomic cosets modulo $nr$ whose coset leaders are in $\Gamma_{(q,n,r)}^{(1)}$ and sizes are equal to $l$, that is,
$N_{l}^{(q,n,r)}=\Big|\big\{C_{i}:\,|C_{i}|=l,i\in Z_{n,r}\big\}\Big|.$
By Lemma 2.1, if $l\nmid\mathrm{ord}_{nr}(q)$, then $N_{l}^{(q,n,r)}=0$.
@@
$N_{l}=N_{l}^{(q,n,r)},\quad l\geq 1.$
We recall two lemmas in [42, Lemmas 4 and 5], which are useful for determining $N_{l}$.
@@
From Lemma 2.3 and using the Möbius inversion formula, we directly get the following formula about the number of elements $i\in Z_{n,r}$ with $l_{i}$ dividing a given integer.
@@ 2.2 Coset leaders in $Z_{n,r}$
For any positive integer $l$, if there exist $q$-cyclotomic cosets with $l$ elements (that is, $N_{l}^{(q,n,r)}>0$) in $Z_{n,r}$, then we list their coset leaders in ascending order, and we denote the $i$-th one by $\delta_{i}^{(q,n,r,l)}$ for $1\leq i\leq N_{l}^{(q,n,r)}$.
@@
$N_{l}=N_{l}^{(q,n,r)},\quad l\geq 1,$
$\delta_{i}^{(l)}=\delta_{i}^{(q,n,r,l)},\quad 1\leq i\leq N_{l}.$
We want to get some estimates on the coset leaders $\delta_{i}^{(l)}$ for $1\leq i\leq N_{l}$. For this, we need a preparation.
@@
Assume that the $i$-th element not disivible by $q$ in the set $S_{a,r}$ is a+(j1)ra+(j-1)r for some integer $j$. Then, we have either $i=j-\lfloor j/q\rfloor-1$, or $i=j-\lfloor j/q\rfloor$. If $i=j-\lfloor j/q\rfloor-1$, then we have $i\leq j-(j/q-(q-1)/q)-1,$ which gives $j\geq\frac{qi}{q-1}+\frac{1}{q-1}.$ If $i=j-\lfloor j/q\rfloor$, then we have $i\leq j-(j/q-(q-1)/q),$ which implies $j\geq\frac{qi}{q-1}-1.$ Hence, noticing that $j$ is an integer, we always have
@@
First,by Lemma 2.2 we have $\frac{nr}{\gcd(nr,q^{l}-1)}\,\Big|\,\delta_{i}^{(l)}.$ For simplicity, we denote $t=nr/\gcd(nr,q^{l}-1)$, then $\delta_{i}^{(l)}$ is a multiple of $t$. Since $r\mid q-1$, we have $t\mid n$.
@@
$S(t)=\{at,at+rt,at+2rt,\ldots,at+jrt\}=\{at,(a+r)t,(a+2r)t,\ldots,(a+jr)t\},$
where $j$ is the largest integer such that $at+jrt\leq 1+(n-1)r$. Hence, for our purpose it suffices to estimate the $i$-th element in $S(t)$ (listed in ascending order) which is not disivible by $q$.
@@
$\begin{split}\delta_{i}^{(l)}\geq\left(a+(\lceil qi/(q-1)\rceil-2)r\right)t\geq\frac{nr(1+(\lceil qi/(q-1)\rceil-2)r)}{\gcd(nr,q^{l}-1)}.\end{split}$
@@
The following lemma is a generalization of [39, Lemma 9] for determining some elements which are not coset leaders.
@@
Moreover, all these $q^{i}+qt+j$ are less than $nr$ and pairwise distinct.
@@
$\begin{split}q^{i}+qt+j<q^{i}+q\cdot\frac{q^{i}-(q^{2}-1)-1}{q^{3}-q}+q-1<2q^{i}\leq q^{i+1}.\end{split}$ (2.1)